### v1

# INSIGHT: Universal Neural Simulator for Analog Circuits Harnessing Autoregressive Transformers

Abstract.
Analog front-end design heavily relies on specialized human expertise and costly trial-and-error simulations, which motivated many prior works on analog design automation. However, efficient and effective exploration of the vast and complex design space remains constrained by the time-consuming nature of CPU-based SPICE simulations, making effective design automation a challenging endeavor. In this paper, we introduce INSIGHT, a GPU-powered, technology-independent, effective universal neural simulator in the analog front-end design automation loop. INSIGHT accurately predicts the performance metrics of analog circuits across various technology nodes, significantly reducing inference time. Notably, its autoregressive capabilities enable INSIGHT to accurately predict simulation-costly critical transient specifications leveraging less expensive performance metric information. The low cost and high fidelity feature make INSIGHT a good substitute for standard simulators in analog front-end optimization frameworks. INSIGHT is compatible with any optimization framework, facilitating enhanced design space exploration for sample efficiency through sophisticated offline learning and adaptation techniques. Our experiments demonstrate that INSIGHT-M, a model-based batch reinforcement learning framework that leverages INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits. To the best of our knowledge, this marks the first use of autoregressive transformers in analog front-end design.
1. Introduction
Analog circuits vary significantly across and within types, each characterized by distinct performance metrics, making analog front-end design challenging. The heavy reliance on human designer expertise and numerous rounds of costly trial-and-error simulations make the process labor-intensive. The consequent prolonged design cycles and increased time to market have motivated extensive prior research towards its automation.
Various knowledge-based and optimization-based approaches have been proposed to automate analog front-end design, addressing circuit topology selection and device sizing. Knowledge-based methods, such as  1, 2, 3, 4, 5, are time-consuming and lack scalability. Optimization-based approaches, on the other hand, can be split into two categories: equation-based and simulation-based. Equation-based methods like  6, 7, 8, 9, 10, 11, 12 are highly inaccurate on modern technologies and lack scalability, while the simulation-based ones rely on expensive simulations. Evolutionary algorithms like Particle swarm optimization (PSO)  13, advanced differential evolution  14, and genetic algorithms  15, 16 require extensive SPICE simulations (sample inefficient). Bayesian optimization  17, 18 suffers from scalability issues arising from computational demands associated with model updates. Deep Reinforcement Learning (DRL) methods like  19, 20, 21, 22, 23, 24 demonstrate robustness but involve numerous simulator interactions. Modern algorithms  25, 26, 27 use black-box optimization and Model-Based Reinforcement Learning (RL) for device sizing. While they demonstrate improved sample efficiency, they still require many costly simulations. Given the vast and complex design space and the intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance. As a result, data-driven approaches have emerged as a promising solution, representing a recent trend in the field 28, 29.
Initially developed for language tasks 30, 31, transformer architectures 32 have been successfully extended to a diverse array of fields that involve recognizing complex patterns in the data, including vision 33, graph theory 34, and reinforcement learning 35, 36. This expansion was driven by the insight that many problems can be modeled as autoregressive sequence generation tasks. Autoregressive approaches significantly enhance the accuracy of output predictions, ensuring coherent, contextually aware outputs, compared to non-autoregressive approaches, with a minor trade-off in inference speed  37. Transformers have outperformed traditional models like convolutional 38, recurrent 39, and graph neural networks 40, demonstrating superior capabilities in handling complex data. Moreover, their ability to develop foundation models that learn generalized features from data, which can then be fine-tuned for specific tasks with exceptional performance, exemplified by GPT-3 30, has set new trends in AI. However, despite their ability to handle varying input and output sequence lengths, their application for building performance prediction models for analog circuits with diverse circuit parameters and specifications lengths remains unexplored.
In this paper, we introduce INSIGHT, a novel GPU-powered technology-independent universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and different technologies. Leveraging transformers’ inherent strengths, such as high adaptability and computational efficiency, INSIGHT consistently demonstrates high R2 scores and minimal test losses. Its autoregressive capabilities facilitate accurate prediction of simulation-costly critical transient specifications leveraging performance data from less expensive simulations. These features position INSIGHT as a cost-effective, high-fidelity alternative to standard circuit simulators in analog front-end optimization frameworks. INSIGHT is highly scalable and can seamlessly integrate into existing optimization workflows, facilitating computationally economical and constraint-free design space exploration. This integration significantly enhances traditional EDA optimization schemes, which often face bottlenecks due to computational constraints arising from the compute-intensive nature of simulations required at each design cycle iteration. Additionally, transformers’ robust transfer learning capabilities facilitate INSIGHT’s smooth adaptation to evolving technology landscapes while minimizing data requirements. We integrate INSIGHT within a Model-based Batch RL framework (INSIGHT-M) for circuit sizing and observe very high sample efficiencies for different circuits, underscoring the transformative capabilities of transformers in Analog Front-end Design Automation. To our knowledge, this marks the first application of transformers in Analog Front-end Design Automation.
We have validated our concept on different circuits covering different circuit types including operational amplifiers, trans-impedance amplifiers, comparators, and level shifters across different technology nodes.
The key contributions of this work include:
  • We introduce a novel approach to performance prediction for analog circuits by conceptualizing it as a sequence modeling problem. Inspired by recent advancements in natural language processing, we adapt transformer architectures, which are pivotal to modern language models, to decode intricate behaviors in circuits, thereby enhancing predictive accuracy and efficiency.
  • We propose INSIGHT, an innovative GPU-accelerated technology-independent universal neural circuit simulator. INSIGHT is designed to predict performance across diverse circuits and parameters, offering a versatile tool that significantly improves prediction speed and accuracy.
  • We have rigorously validated our approach across six different circuits across multiple technology nodes. INSIGHT consistently achieves high R2 scores and minimal loss scores across circuits and technology nodes. Its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information, thus demonstrating its capability as a high-fidelity substitute for standard simulators. By embedding INSIGHT within a Model-based Batch RL framework, INSIGHT-M, for circuit sizing optimization, we achieved exceptional sample efficiencies, with at least 50×50$\times$50 × improvement across circuits, thus underscoring INSIGHT’s potential to revolutionize diverse optimization frameworks.
  • To the best of our knowledge, this is the first application of transformer models in analog front-end design automation, paving the way for leveraging recent advancements in language models to enhance circuit design and optimization.
The rest of the paper is organized as follows. Section 2 describes the background and motivation; Section 3 explains the detailed implementation; Section 4 demonstrates the results; Section 5 concludes the paper.
2. Preliminaries
In this section, we delve into the fundamentals of analog circuit physics (Section  2.1) and explore transformer architectures (Section  2.2), discussing our rationale for choosing them. We also highlight the definition of a key design quality metric in Section  2.3, called Figure of Merit (FOM), which will guide our sizing flow.
2.1. Analog Circuit Physics
In analog circuits, performance metrics and parameters are highly interdependent. By leveraging these interdependencies, we can build effective prediction models. For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, including unity-gain bandwidth (UGBW) and phase margin (PM). Similarly, combining all these metrics alongside parameter information provides crucial insights into the circuit’s potential transient behavior. Thus, integrating partial performance information alongside parameter information can significantly enhance prediction accuracies for the other performance specifications, including critical transient specifications such as slew rate and settling time, which are typically costly and slow to simulate. This insight motivates the adoption of autoregressive models for precise performance prediction. These models leverage partial performance information along with parameter information to provide a richer context for predicting the remaining metrics, aiming to build a high-fidelity substitute for standard simulators. Furthermore, given the transferable nature of circuit physics across technology nodes, understanding performance interdependencies within one node provides valuable insights applicable to other nodes. Consequently, a model exhibiting high predictive accuracy for circuit performance on one technology node will likely achieve similar success on subsequent nodes and be easily transferable across advanced technology nodes. These insights, combined with the adaptability and transferability of transformer architectures, motivate the use of autoregressive transformers to build a universal neural simulator that would work efficiently across technologies.
2.2. Transformer Architectures
Transformers  32 constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond. These models, foundational for various state-of-the-art language models, stand out due to their self-attention mechanism, effectively capturing dependencies between tokens in a sequence, regardless of distance. The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors. These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors. Additionally, unlike traditional RNNs 41, transformers process data in parallel, significantly improving efficiency and scalability.
While the original Transformer architecture 32 consists of stacked layers of encoders and decoders, decoder-only models like GPT 42 have recently demonstrated notable success in sequence data generation. Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks. Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence. This architectural choice is optimized for autoregressive tasks, where the model iteratively predicts the subsequent token by leveraging all previously generated tokens along with the input sequence as context.
From a probabilistic perspective, in the context of autoregressive generation, transformers model the joint probability of a sequence of tokens ($x_{1}$,$x_{2}$,…,$x_{N}$) by decomposing it into a product of conditional probabilities:
(1) $P(x_{1},x_{2},\dots,x_{N})=\prod_{i=1}^{N}P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$
Each term $P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$, which quantifies the likelihood of observing the token $x_{i}$, given the sequence of all preceding tokens ($x_{1},x_{2},\dots,x_{i-1}$), is estimated using the self-attention mechanism. The mechanism’s ability to capture long-range dependencies and contextual relationships within data, along with its adaptability to variable sequence lengths, makes Transformers ideal for handling complex sequences with widely varying input sizes. Drawing inspiration from this probabilistic formulation, we aim to enhance analog circuit performance prediction by strategically integrating partial performance data with parameter information, providing richer context as detailed in the following section.
2.3. Figure of Merit
Analog circuit sizing involves determining the optimal device sizes to meet specified performance criteria. Formally, given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  25, 43, 26, to evaluate quality of a design. This is of the following form:
(2) $\text{FoM}(\mathbf{x})=w_{0}\cdot f_{0}(\mathbf{x})+\sum_{i=1}^{m}\min\left(1,\max(0,w_{i}\cdot f_{i}(\mathbf{x}))\right)$
where $w_{i}$ is the weighting factor, a $\max(\cdot)$ function is used for equating designs after constraints are met, and $\min(\cdot)$ is used to prevent a single constraint violation from dominating the FoM value.
3. Algorithms
In this section, we introduce our proposed frameworks. We begin by detailing the construction of our proposed universal neural simulator for analog circuits, INSIGHT. Figure 1 illustrates how INSIGHT conceptualizes the prediction of performance metrics as an autoregressive sequence generation task, with Section 3.1 providing the probabilistic rationale for this approach. Section 3.2 then offers a detailed discussion of our analog circuit sizing optimization algorithm leveraging INSIGHT. As illustrated in Figure 2, the analog circuit sizing optimization algorithm for a given topology consists of two primary phases. The initial phase, discussed in Section 3.2.1, consists of a one-time effort for simulation-based performance dataset generation and training of INSIGHT. The second phase, discussed in Section 3.3, involves the integration of the one-time trained INSIGHT into a Model-Based Batch RL framework, termed INSIGHT-M, to perform automated analog sizing optimization for previously unseen user-input target specifications.
3.1. INSIGHT
Formally, for a given analog circuit topology, given a sequence of design parameters ($x_{1}$ to $x_{N}$), such as width-to-length ratios (W/L), our goal is to predict the corresponding performance specifications ($y_{1}$ to $y_{M}$) accurately. This predictive task is mathematically framed as maximizing the conditional probability $P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})$, which can be expressed as follows:
(3) $\begin{split}P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})=\\ \prod_{i=1}^{M}P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})\end{split}$
This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
(4) $\text{NLL}=-\sum_{i=1}^{M}\log P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})$
This formulation mirrors the contextual approach used in autoregressive sequence generation tasks, such as those involving decoder-based transformers designed for enhanced sequence analysis. In analog circuits, where performance metrics are inherently interdependent, this formulation attempts to leverage these intricate relationships to predict the performance metrics for previously unseen input parameters accurately. Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics. The self-attention mechanism within the Transformer model enables it to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$. This capability is crucial for uncovering insights into the physics of the circuit, as it allows the model to learn and utilize the dependencies of one performance metric on another along with parameter information. Table 1 presents INSIGHT’s architecture details.
Figure 1. Autoregressive circuit performance prediction in INSIGHT
3.2. Analog Circuit Sizing Optimization
Figure 2. Automatic Analog Sizing Framework
Figure 2 showcases our sizing optimization algorithm. It comprises of two main sections: INSIGHT Pre-Training Phase (Section  3.2.1) and INSIGHT-M (Section  3.3). The core of RL in this paper is PPO 44. We basically use the hyper-parameters therein. Note that other RL algorithms can be employed with INSIGHT.
Table 1. INSIGHT Architecture
INSIGHT Parameter Value
  Network Architecture Decoder-Based Transformer
  Dim. of a model 76
  # of Heads 4
  # of Decoder Layers 3
  Activation GeLU
Table 2. Neural Simulator Architecture in CRONuS
FC-Ensemble Parameter Value
 Network Architecture Fully-Connected
# of Hidden Units 200-200-200-200-200
# of Ensemble Networks 7
Activation ReLU
3.2.1   INSIGHT Pre-Training
As depicted in the left part of Figure 2, the pre-training phase involves a one-time effort to generate a simulation-based performance dataset, followed by a one-time training of our proposed neural simulator, INSIGHT. The data collection involves gathering a random dataset of Gaussian-sampled design parameters and their corresponding performance metrics derived from SPICE simulations. To address computational demands, the initial dataset size is defined as a hyperparameter, allowing for adjustments tailored to specific use cases.
This flexibility helps balance the resource usage during the initial data collection with the subsequent sample efficiency following a user query. This process is followed by the one-time pre-training of our proposed neural simulator, INSIGHT. By employing transformer technology, INSIGHT effectively utilizes transfer learning, thereby reducing dataset requirements across different technologies.
3.3. INSIGHT-M
The pre-trained INSIGHT is subsequently integrated into a Model-Based Batch RL framework for circuit sizing, referred to as INSIGHT-M. As shown in the right part of Figure 2, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance target specifications, including previously unseen ones. In our experiments, we utilized a Proximal Policy Optimization (PPO)-based RL agent 44, a widely used approach in Reinforcement Learning. PPO has been extensively utilized in analog design optimization, as in  45, 24. PPO operates by executing multiple trajectory rollouts in parallel, followed by their subsequent aggregation to update the policy under certain constraints. Although PPO’s use of multiple trajectories under safe constraints enhances the robustness and safety of the learning process, it tends to compromise sample efficiency. The INSIGHT-M framework addresses this drawback by enabling the RL agents to be pre-trained within the close-to-real synthetic neural simulator environment, INSIGHT, until they achieve satisfactory performance. Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy. This process involves periodic performance evaluations in the real environment. If the agents perform adequately, their deployment persists. If not, new real-world data is acquired to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated. This iterative refinement process optimizes the balance between safety, robustness, and efficiency.
Central to the efficiency of INSIGHT-M is the training of a robust RL agent within a synthetic environment that closely mimics real-world conditions. The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency. Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, enhances the agent’s adaptability to real-world environments. Furthermore, the framework’s capability for iterative refinement effectively balances exploration and exploitation, achieving high sample efficiency.
4. Experiments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e) (f)
Figure 3. Benchmark Circuits: (a) Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, (b) Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, (c) Two-Stage Transimpedance Amplifier, (d) Three-Stage Transimpedance Amplifier, (e) Comparator, (f) Level Shifter.
Table 3. Performance Comparison Across Different Circuits: INSIGHT vs Neural Simulator in CRONuS
Circuits (45nm) R2 Scores
Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¿0.99 0.65 ¿0.99 0.62
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¿0.99 0.73 ¿0.99 0.77
Two-Stage Transimpedance Amplifier ¿0.99 0.81 ¿0.99 0.78
Three-Stage Transimpedance Amplifier ¿0.99 0.91 ¿0.99 0.88
Comparator 0.70 0.60 0.97 0.61
Level Shifter ¿0.99 ¡0.00 0.91 0.07
Circuits (45nm) MSE
Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¡0.001 0.01 ¡0.001 0.01
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¡0.001 0.05 ¡0.001 0.04
Two-Stage Transimpedance Amplifier ¡0.001 0.05 ¡0.001 0.05
Three-Stage Transimpedance Amplifier ¡0.001 0.01 ¡0.001 0.01
Comparator 0.010 0.01 ¡0.001 0.01
Level Shifter ¡0.001 0.03 ¡0.001 0.05
Table 4. INSIGHT Performance Across Different Technologies
Technology 45nm 130nm 180nm
Circuit Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Two-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Three-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Table 5. Performance Comparison of INSIGHT-M (with PPO) vs AutoCkt for 45nm case
Algorithm INSIGHT-M (with PPO) AutoCkt
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1,500 steps 14,100 steps
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1,500 steps 9,300 steps
Two-Stage Transimpedance Amplifier 300 steps 11,100 steps
Three-Stage Transimpedance Amplifier 600 steps 8,700 steps
Comparator 1,200 steps 8,400 steps
4.1. Benchmark Tasks and Algorithms
In this paper, we utilized a GPU machine with x8 NVIDIA P100 GPUs under Intel(R) Xeon(R) CPU E5-2698 V4 2.220GHz for training and inference. We used the following benchmark circuits to evaluate INSIGHT and to validate our approach: Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, Two-Stage Transimpedance Amplifier, Three-Stage Transimpedance Amplifier, Comparator, and Level Shifter. Figure 3 illustrates the detailed topology of the benchmark circuits. Table 6 presents the performance specifications for each circuit we focussed on for this paper. We have validated our concept using BSIM 45nm, Skywater 130nm, BSIM 180nm, and FreePDK 45nm technology nodes.
4.2. INSIGHT Performance
4.2.1   Prediction Accuracy
Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  27, which serves as the baseline for our comparison with INSIGHT. Table 3 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on the BSIM 45nm technology node. As demonstrated in the table, our analysis for the comparator and level shifter cases involved a combination of DC and simulation-costly transient specifications. Specifically for the level shifter, leveraging its autoregressive capabilities, INSIGHT enhances the predictability of simulation-intensive transient specifications such as Avg. Delay and Delay Balance by efficiently utilizing economical metric information like DC Power and Ratio. The table further details INSIGHT’s performance across various training and testing data sizes, indicating robust performance for dataset sizes close to 1500-1800. Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times. In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases. Table 4 validates INSIGHT’s effectiveness across 45nm, 130nm, and 180nm technology nodes, reinforcing its universal applicability. Our experiments showed that simple transfer learning or knowledge-based domain adaptation techniques can reduce initial dataset requirements.
4.2.2   Inference Speed
We evaluated INSIGHT’s performance prediction inference time across our benchmark circuits. When simultaneously processing a batch of 1,000 combinations of input parameters, the average inference time across circuits was $115.21\pm 12.01\mu s$. This result was derived from 100 trials per circuit on our setup.
4.3. INSIGHT-M Performance
We evaluate the performance of INSIGHT-M against AutoCkt 24, an established online learning-based sizing algorithm based on Proximal Policy Optimization (PPO), using default hyperparameter settings. Table 5 highlights the performance comparison. Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance. For these results, the INSIGHT model used within INSIGHT-M was pre-trained on a minimal initial dataset of 300 samples, showcasing the efficacy of our neural simulator INSIGHT in optimization frameworks.
Table 6. Performance Specifications for each circuit
Circuit Performance Specifications Width (Capacitor) Range ($\mu$m ($pF$))
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair $I_{Q}$ (mA) DC Gain (dB) UGBW (MHz) Phase Margin (degrees) [1.0e-1, 3.0e1]
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair [1.0e-1, 3.0e1]
Two-Stage Transimpedance Amplifier [1.0e-1, 1.0e2]
Three-Stage Transimpedance Amplifier [1.0e-2, 3.0e1]
Comparator DC Power (W) Avg. Delay (sec) [1.0e-1, 3.0e1]
Level Shifter DC Power (W) Ratio Avg. Delay (sec) Delay Balance (sec) [4.6e-1, 1.0e2]
5. Conclusion
This paper proposes INSIGHT, a novel, GPU-powered, technology-independent, effective universal neural simulator in the analog front-end automation loop. Leveraging self-attention to understand underlying inherent parameter-performance interdependencies, INSIGHT accurately predicts performance across analog circuits and their corresponding metrics over various technology nodes, achieving inference times within microseconds. Notably, its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information. Due to its low cost and high fidelity, INSIGHT serves as an effective substitute for standard simulators in analog front-end optimization frameworks, facilitating effective and efficient design space exploration and improved sample efficiency. Our Model-Based Batch RL framework, INSIGHT-M, which utilizes INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits. We rigorously validated our approach on diverse circuit types across technology nodes. To our knowledge, this marks the first application of autoregressive transformers in analog front-end design automation. We plan to extend this technology to address more complex analog design challenges, including mismatch analysis and design migration. We aim to enhance our adaptation techniques to mitigate synthetic to real-world adaptation errors, thereby strengthening the robustness and reliability of our models under dynamic conditions. This innovative application of autoregressive transformers opens new avenues for leveraging advanced NLP techniques, including grammatical models and language-based inversion techniques in circuit modeling and optimization, inspiring further innovation in the field.

### v2

@@
 Analog front-end design heavily relies on specialized human expertise and costly trial-and-error simulations, which motivated many prior works on analog design automation.
-However, efficient and effective exploration of the vast and complex design space remains constrained by the time-consuming nature of CPU-based SPICE simulations, making effective design automation a challenging endeavor.
-In this paper, we introduce INSIGHT, a GPU-powered, technology-independent, effective universal neural simulator in the analog front-end design automation loop.
-INSIGHT accurately predicts the performance metrics of analog circuits across various technology nodes, significantly reducing inference time.
+However, efficient and effective exploration of the vast and complex design space remains constrained by the time-consuming nature of SPICE simulations, making effective design automation a challenging endeavor.
+In this paper, we introduce INSIGHT, a GPU-powered, technology-agnostic, effective universal neural simulator in the analog front-end design automation loop.
+INSIGHT accurately predicts the performance metrics of analog circuits across various technologies with just a few microseconds of inference time.
 Notably, its autoregressive capabilities enable INSIGHT to accurately predict simulation-costly critical transient specifications leveraging less expensive performance metric information.
@@
 INSIGHT is compatible with any optimization framework, facilitating enhanced design space exploration for sample efficiency through sophisticated offline learning and adaptation techniques.
-Our experiments demonstrate that INSIGHT-M, a model-based batch reinforcement learning framework that leverages INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits.
-To the best of our knowledge, this marks the first use of autoregressive transformers in analog front-end design.
+Our experiments demonstrate that INSIGHT-M, a model-based batch reinforcement learning sizing framework with INSIGHT as the accurate surrogate, only requires $<$ 20 real-time simulations with 100-1000x lower simulation costs and significant speedup over existing sizing methods.
 1. Introduction
 Analog circuits vary significantly across and within types, each characterized by distinct performance metrics, making analog front-end design challenging.
-The heavy reliance on human designer expertise and numerous rounds of costly trial-and-error simulations make the process labor-intensive.
-The consequent prolonged design cycles and increased time to market have motivated extensive prior research towards its automation.
-Various knowledge-based and optimization-based approaches have been proposed to automate analog front-end design, addressing circuit topology selection and device sizing.
-Knowledge-based methods, such as  1, 2, 3, 4, 5, are time-consuming and lack scalability.
-Optimization-based approaches, on the other hand, can be split into two categories: equation-based and simulation-based.
-Equation-based methods like  6, 7, 8, 9, 10, 11, 12 are highly inaccurate on modern technologies and lack scalability, while the simulation-based ones rely on expensive simulations.
-Evolutionary algorithms like Particle swarm optimization (PSO)  13, advanced differential evolution  14, and genetic algorithms  15, 16 require extensive SPICE simulations (sample inefficient).
-Bayesian optimization  17, 18 suffers from scalability issues arising from computational demands associated with model updates.
-Deep Reinforcement Learning (DRL) methods like  19, 20, 21, 22, 23, 24 demonstrate robustness but involve numerous simulator interactions.
-Modern algorithms  25, 26, 27 use black-box optimization and Model-Based Reinforcement Learning (RL) for device sizing.
-While they demonstrate improved sample efficiency, they still require many costly simulations.
-Given the vast and complex design space and the intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance.
-As a result, data-driven approaches have emerged as a promising solution, representing a recent trend in the field 28, 29.
-Initially developed for language tasks 30, 31, transformer architectures 32 have been successfully extended to a diverse array of fields that involve recognizing complex patterns in the data, including vision 33, graph theory 34, and reinforcement learning 35, 36.
-This expansion was driven by the insight that many problems can be modeled as autoregressive sequence generation tasks.
-Autoregressive approaches significantly enhance the accuracy of output predictions, ensuring coherent, contextually aware outputs, compared to non-autoregressive approaches, with a minor trade-off in inference speed  37.
-Transformers have outperformed traditional models like convolutional 38, recurrent 39, and graph neural networks 40, demonstrating superior capabilities in handling complex data.
-Moreover, their ability to develop foundation models that learn generalized features from data, which can then be fine-tuned for specific tasks with exceptional performance, exemplified by GPT-3 30, has set new trends in AI.
-However, despite their ability to handle varying input and output sequence lengths, their application for building performance prediction models for analog circuits with diverse circuit parameters and specifications lengths remains unexplored.
-In this paper, we introduce INSIGHT, a novel GPU-powered technology-independent universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and different technologies.
-Leveraging transformers’ inherent strengths, such as high adaptability and computational efficiency, INSIGHT consistently demonstrates high R2 scores and minimal test losses.
+This complexity necessitates a labor-intensive process reliant on designer expertise and costly trial-and-error simulations.
+The consequent prolonged design cycles and increased time to market have motivated extensive research towards its automation.
+Various knowledge-based and optimization-based approaches, including equation-based and simulation-based methods, have been proposed for analog design automation.
+While knowledge-based methods, such as  1, 2, 3, 4, are time-consuming and lack scalability, equation-based methods, like  5, 6, 7, 8, are also highly inaccurate on modern technologies.
+Conversely, simulation-based methods, while effective, depend on costly real-time simulations.
+Among these, evolutionary algorithms like particle swarm optimization  9, advanced differential evolution  10, and genetic algorithms  11, 12 require extensive SPICE simulations (sample inefficient).
+Deep Reinforcement Learning (RL) methods like  13, 14, 15, 16, 17, 18, 19, 20 while robust, also involve frequent simulator interactions.
+Meanwhile, Bayesian Optimization  21, 22 suffers from scalability issues due to computational demands associated with model updates.
+Modern algorithms  23, 24, 25 use black-box optimization and model-based RL to enhance sample efficiency, yet they still require numerous costly simulations.
+Given the vast and complex design space and intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance.
+This heavy reliance on expensive real-time simulations motivates us to find an efficient and effective surrogate for the standard SPICE simulator to expedite analog design optimization tasks.
+The complexity of parameter-performance relationships in analog circuits renders simple DNNs inadequate, necessitating bigger, more complex models and extensive, costly EDA data while still resulting in suboptimal accuracies 25.
+Given the high interdependence of performance metrics inherent to analog circuits, we propose to formulate the analog circuit performance prediction problem as an autoregressive generation task, i.e., predicting each remaining simulation-costly performance specification by leveraging provided design parameters and partial performance information (provided or previously-generated) in a sequential manner.
+The autoregressive problem setup enforces the model to automatically reason and capture the inherent interdependency in the challenging analog performance prediction task, easing the model learning complexity, therefore, providing much better prediction quality than vanilla DNN setup 25.
+In this paper, we introduce INSIGHT, a novel, technology-agnostic, universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and technologies autoregressively, consistently demonstrating R2 scores as high as 0.99 and minimal test losses at microsecond inference times.
 Its autoregressive capabilities facilitate accurate prediction of simulation-costly critical transient specifications leveraging performance data from less expensive simulations.
-These features position INSIGHT as a cost-effective, high-fidelity alternative to standard circuit simulators in analog front-end optimization frameworks.
-INSIGHT is highly scalable and can seamlessly integrate into existing optimization workflows, facilitating computationally economical and constraint-free design space exploration.
-This integration significantly enhances traditional EDA optimization schemes, which often face bottlenecks due to computational constraints arising from the compute-intensive nature of simulations required at each design cycle iteration.
-Additionally, transformers’ robust transfer learning capabilities facilitate INSIGHT’s smooth adaptation to evolving technology landscapes while minimizing data requirements.
-We integrate INSIGHT within a Model-based Batch RL framework (INSIGHT-M) for circuit sizing and observe very high sample efficiencies for different circuits, underscoring the transformative capabilities of transformers in Analog Front-end Design Automation.
-To our knowledge, this marks the first application of transformers in Analog Front-end Design Automation.
-We have validated our concept on different circuits covering different circuit types including operational amplifiers, trans-impedance amplifiers, comparators, and level shifters across different technology nodes.
+These features position INSIGHT as a cost-effective and high-fidelity alternative to standard circuit simulators in analog front-end optimization frameworks.
+INSIGHT is highly scalable and can integrate into existing EDA optimization workflows to enhance efficiency.
+To showcase the effectiveness of INSIGHT as a surrogate model, we integrate INSIGHT within a model-based batch RL framework (INSIGHT-M) for circuit sizing and observe very high sample efficiencies for different circuits, covering several circuit types, like operational amplifiers, transimpedance amplifiers, comparators, and level shifters, across various technologies.
 The key contributions of this work include:
-  • We introduce a novel approach to performance prediction for analog circuits by conceptualizing it as a sequence modeling problem.
-Inspired by recent advancements in natural language processing, we adapt transformer architectures, which are pivotal to modern language models, to decode intricate behaviors in circuits, thereby enhancing predictive accuracy and efficiency.
-  • We propose INSIGHT, an innovative GPU-accelerated technology-independent universal neural circuit simulator.
-INSIGHT is designed to predict performance across diverse circuits and parameters, offering a versatile tool that significantly improves prediction speed and accuracy.
-  • We have rigorously validated our approach across six different circuits across multiple technology nodes.
-INSIGHT consistently achieves high R2 scores and minimal loss scores across circuits and technology nodes.
-Its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information, thus demonstrating its capability as a high-fidelity substitute for standard simulators.
-By embedding INSIGHT within a Model-based Batch RL framework, INSIGHT-M, for circuit sizing optimization, we achieved exceptional sample efficiencies, with at least 50×50$\times$50 × improvement across circuits, thus underscoring INSIGHT’s potential to revolutionize diverse optimization frameworks.
-  • To the best of our knowledge, this is the first application of transformer models in analog front-end design automation, paving the way for leveraging recent advancements in language models to enhance circuit design and optimization.
-The rest of the paper is organized as follows.
-Section 2 describes the background and motivation; Section 3 explains the detailed implementation; Section 4 demonstrates the results; Section 5 concludes the paper.
+  • We introduce a novel problem formulation for analog circuit performance prediction by conceptualizing it as an autoregressive sequence modeling problem, which significantly eases the learning complexity of the neural model.
+  • We build INSIGHT, a GPU-accelerated, technology-agnostic universal neural simulator based on a decoder-only Transformer architecture to solve the analog circuit performance prediction problem.
+INSIGHT enables accurate prediction of simulation-costly critical transient specifications using less expensive performance metrics and and demonstrates its potential as a low-cost, high-fidelity surrogate model in the analog design automation loop.
+  • We have rigorously validated our INSIGHT across various circuit and performance metric types across multiple technologies.
+We also demonstrate INSIGHT-M, which leverages INSIGHT as the performance predictor for the device sizing problem.
+INSIGHT-M achieved exceptional sample efficiencies on various circuit types, requiring $<$ 20 real-time simulations compared to existing sizing methods, which require 100-1000x more simulations.
+  • To the best of our knowledge, this marks the first successful application of a universal neural simulator framework leveraging autoregressive Transformers for analog design automation.
 2. Preliminaries
-In this section, we delve into the fundamentals of analog circuit physics (Section  2.1) and explore transformer architectures (Section  2.2), discussing our rationale for choosing them.
-We also highlight the definition of a key design quality metric in Section  2.3, called Figure of Merit (FOM), which will guide our sizing flow.
-2.1. Analog Circuit Physics
+In this section, we briefly discuss some insights into analog circuit basics/metrics (Section  2.1) and Transformer architectures (Section  2.2), discussing our rationale for choosing them.
+Additionally, we define a key design quality metric, the Figure of Merit (FoM), which will guide our sizing flow (Section  2.3).
+2.1. Analog Circuit Basics and Metrics
 In analog circuits, performance metrics and parameters are highly interdependent.
 By leveraging these interdependencies, we can build effective prediction models.
-For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, including unity-gain bandwidth (UGBW) and phase margin (PM).
-Similarly, combining all these metrics alongside parameter information provides crucial insights into the circuit’s potential transient behavior.
-Thus, integrating partial performance information alongside parameter information can significantly enhance prediction accuracies for the other performance specifications, including critical transient specifications such as slew rate and settling time, which are typically costly and slow to simulate.
+For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, like unity-gain bandwidth (UGBW) and phase margin (PM).
+Similarly, integrating all static metrics provides crucial insights into the circuit’s potential transient behavior.
+Thus, leveraging less expensive performance metric information can significantly enhance prediction accuracies for critical, simulation-costly, transient specifications such as slew rate and settling time.
 This insight motivates the adoption of autoregressive models for precise performance prediction.
-These models leverage partial performance information along with parameter information to provide a richer context for predicting the remaining metrics, aiming to build a high-fidelity substitute for standard simulators.
-Furthermore, given the transferable nature of circuit physics across technology nodes, understanding performance interdependencies within one node provides valuable insights applicable to other nodes.
-Consequently, a model exhibiting high predictive accuracy for circuit performance on one technology node will likely achieve similar success on subsequent nodes and be easily transferable across advanced technology nodes.
-These insights, combined with the adaptability and transferability of transformer architectures, motivate the use of autoregressive transformers to build a universal neural simulator that would work efficiently across technologies.
+Furthermore, given the transferable nature of circuit physics across technologies for a given topology, understanding performance interdependencies within one node provides valuable insights applicable to other nodes.
+These insights, combined with the adaptability and transferability of Transformer architectures, motivate the adoption of autoregressive Transformers to build a universal neural simulator that would work efficiently across technologies.
 2.2. Transformer Architectures
-Transformers  32 constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond.
-These models, foundational for various state-of-the-art language models, stand out due to their self-attention mechanism, effectively capturing dependencies between tokens in a sequence, regardless of distance.
+Transformers constitute a powerful class of generative deep learning architectures that have revolutionized sequence modeling tasks in natural language processing (NLP) and beyond.
+These models stand out due to their self-attention mechanism, which effectively captures dependencies between tokens in a sequence regardless of distance.
 The self-attention mechanism evaluates dependencies within a sequence by computing a weighted sum of value vectors.
 These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors.
-Additionally, unlike traditional RNNs 41, transformers process data in parallel, significantly improving efficiency and scalability.
-While the original Transformer architecture 32 consists of stacked layers of encoders and decoders, decoder-only models like GPT 42 have recently demonstrated notable success in sequence data generation.
+Unlike traditional RNNs 26, Transformers process data in parallel, enhancing efficiency and scalability.
+While the original Transformer architecture  27 consists of stacked layers of encoders and decoders, decoder-only models like GPT 28 have recently demonstrated notable success in sequence data generation.
 Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks.
-Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence.
-This architectural choice is optimized for autoregressive tasks, where the model iteratively predicts the subsequent token by leveraging all previously generated tokens along with the input sequence as context.
-From a probabilistic perspective, in the context of autoregressive generation, transformers model the joint probability of a sequence of tokens ($x_{1}$,$x_{2}$,…,$x_{N}$) by decomposing it into a product of conditional probabilities:
-(1) $P(x_{1},x_{2},\dots,x_{N})=\prod_{i=1}^{N}P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$
-Each term $P(x_{i}|x_{1},x_{2},\dots,x_{i-1})$, which quantifies the likelihood of observing the token $x_{i}$, given the sequence of all preceding tokens ($x_{1},x_{2},\dots,x_{i-1}$), is estimated using the self-attention mechanism.
-The mechanism’s ability to capture long-range dependencies and contextual relationships within data, along with its adaptability to variable sequence lengths, makes Transformers ideal for handling complex sequences with widely varying input sizes.
-Drawing inspiration from this probabilistic formulation, we aim to enhance analog circuit performance prediction by strategically integrating partial performance data with parameter information, providing richer context as detailed in the following section.
 2.3. Figure of Merit
-Analog circuit sizing involves determining the optimal device sizes to meet specified performance criteria.
-Formally, given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  25, 43, 26, to evaluate quality of a design.
+Given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  23, 29, 24, to evaluate quality of a design.
 This is of the following form:
@@
 where $w_{i}$ is the weighting factor, a $\max(\cdot)$ function is used for equating designs after constraints are met, and $\min(\cdot)$ is used to prevent a single constraint violation from dominating the FoM value.
-3. Algorithms
+This formulation will guide our sizing workflow.
+3. INSIGHT Framework and Algorithms
 In this section, we introduce our proposed frameworks.
-We begin by detailing the construction of our proposed universal neural simulator for analog circuits, INSIGHT.
-Figure 1 illustrates how INSIGHT conceptualizes the prediction of performance metrics as an autoregressive sequence generation task, with Section 3.1 providing the probabilistic rationale for this approach.
-Section 3.2 then offers a detailed discussion of our analog circuit sizing optimization algorithm leveraging INSIGHT.
-As illustrated in Figure 2, the analog circuit sizing optimization algorithm for a given topology consists of two primary phases.
-The initial phase, discussed in Section 3.2.1, consists of a one-time effort for simulation-based performance dataset generation and training of INSIGHT.
-The second phase, discussed in Section 3.3, involves the integration of the one-time trained INSIGHT into a Model-Based Batch RL framework, termed INSIGHT-M, to perform automated analog sizing optimization for previously unseen user-input target specifications.
-3.1. INSIGHT
-Formally, for a given analog circuit topology, given a sequence of design parameters ($x_{1}$ to $x_{N}$), such as width-to-length ratios (W/L), our goal is to predict the corresponding performance specifications ($y_{1}$ to $y_{M}$) accurately.
-This predictive task is mathematically framed as maximizing the conditional probability $P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})$, which can be expressed as follows:
+We begin by detailing the construction of INSIGHT, our proposed universal neural simulator for analog circuits, along with the probabilistic rationale for this approach, in Section 3.1.
+Section  3.2 then details our analog circuit sizing optimization framework for a given topology (Figure 2).
+3.1. INSIGHT for Modeling
+In INSIGHT, we are driven by a critical inquiry: Can we develop an effective and efficient surrogate model that utilizes circuit parameters and inexpensive partial performance metrics information (if available) to predict remaining simulation-costly performance numbers accurately?
+This question is particularly pertinent as current analog front-end design automation heavily depends on time-intensive simulations like SPICE, incurring a strong demand to ease the reliance.
+The problem can be formally represented as follows: For a given analog circuit topology, given a sequence of design parameters ($x_{1}$ to $x_{N}$), such as width-to-length ratios (W/L), our goal is to predict the corresponding performance specifications ($y_{1}$ to $y_{M}$) accurately.
+This predictive task is mathematically framed as maximizing the conditional probability $P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})$, which can be decomposed into a product of conditional probabilities as follows:
 (2) $\begin{split}P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})=\\ \prod_{i=1}^{M}P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})\end{split}$
+Each term $P(y_{i}|x_{1},x_{2},\dots,x_{N},)$, which quantifies the likelihood of observing the token $y_{i}$, given the design parameters and the predicted performance specifications.
 This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
 (3) $\text{NLL}=-\sum_{i=1}^{M}\log P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})$
-This formulation mirrors the contextual approach used in autoregressive sequence generation tasks, such as those involving decoder-based transformers designed for enhanced sequence analysis.
-In analog circuits, where performance metrics are inherently interdependent, this formulation attempts to leverage these intricate relationships to predict the performance metrics for previously unseen input parameters accurately.
+This formulation connects to the contextual approach in autoregressive sequence generation tasks involving decoder-based Transformers like GPT 28 and motivates us to use a decoder-style Transformer model for INSIGHT.
 Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics.
-The self-attention mechanism within the Transformer model enables it to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$.
-This capability is crucial for uncovering insights into the physics of the circuit, as it allows the model to learn and utilize the dependencies of one performance metric on another along with parameter information.
+We utilize the self-attention mechanism to enable the model to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$.
+Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence.
+The autoregressive feature ensures the model learns to utilize dependencies of one performance metric on another, along with the parameter information, to predict the performance metrics accurately.
 Table 1 presents INSIGHT’s architecture details.
-Figure 1. Autoregressive circuit performance prediction in INSIGHT
-3.2. Analog Circuit Sizing Optimization
+Figure 1. Autoregressive Performance Prediction in INSIGHT
 Figure 2. Automatic Analog Sizing Framework
-Figure 2 showcases our sizing optimization algorithm.
-It comprises of two main sections: INSIGHT Pre-Training Phase (Section  3.2.1) and INSIGHT-M (Section  3.3).
-The core of RL in this paper is PPO 44.
-We basically use the hyper-parameters therein.
-Note that other RL algorithms can be employed with INSIGHT.
 Table 1. INSIGHT Architecture
@@
   Network Architecture Decoder-Based Transformer
-  Dim. of a model 76
+  Dim. of model 76
   # of Heads 4
   # of Decoder Layers 3
-  Activation GeLU
+  Activation GeLU 30
+  Learning rate 0.001
+with cosine annealing scheduler  31
 Table 2. Neural Simulator Architecture in CRONuS
@@
 # of Ensemble Networks 7
-Activation ReLU
+Activation ReLU 32
+  Learning rate 0.001
+Although our analog performance prediction task is analogous to the sequence generation tasks, we would like to emphasize that we need to predict real values instead of generating a sequence of tokens, such as a word or a position in an image, from a set of finite elements, or generate scalar real values from sequential contexts, as shown in previous Transformer works.
+Generating a sequence of real values based on Transformer architecture is unclear  33, and thus, suitable foundation models for predicting circuit parameters do not exist.
+Therefore, it is essential to evaluate the feasibility of Transformer architecture in generating a sequence of circuit outputs to confirm the potential for constructing circuit foundation models.
+Accordingly, building on the existing decoder-Transformer architecture 28, 34, we show that our proposed architecture accelerates inference and training to demonstrate the feasibility of a unified framework like a foundation model.
+This streamlined approach enhances processing efficiency, reduces data requirements, and makes the model more practical for various applications.
+By focusing on simplicity, we improve performance and usability while maintaining robust predictions.
+3.2. INSIGHT for Analog Sizing
+Figure 2 showcases our sizing optimization algorithm. It comprises of two main sections: INSIGHT Pre-Training Phase (Section  3.2.1) and INSIGHT for Optimization (Section  3.2.2).
 3.2.1   INSIGHT Pre-Training
@@
 This flexibility helps balance the resource usage during the initial data collection with the subsequent sample efficiency following a user query.
-This process is followed by the one-time pre-training of our proposed neural simulator, INSIGHT.
-By employing transformer technology, INSIGHT effectively utilizes transfer learning, thereby reducing dataset requirements across different technologies.
-3.3. INSIGHT-M
+Additionally, INSIGHT can leverage transfer learning to reduce dataset requirements across technologies while maintaining high performance.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(a) (b) (c) (d) (e) (f)
+Figure 3.
+Benchmark Circuits: (a) 2-Stage Miller-Compensated OPAMP with NMOS Input Pair, (b) 2-Stage Miller-Compensated OPAMP with PMOS Input Pair, (c) 2-Stage TIA, (d) 3-Stage TIA, (e) Comparator, (f) Level Shifter.
+Table 3. Performance Metrics for each circuit
+Circuits Performance Metrics
+2-Stage Miller-Compensated OPAMP with NMOS Input Pair $I_{Q}$ (mA) DC Gain (dB) UGBW (MHz) Phase Margin (${}^{\circ})$
+2-Stage Miller-Compensated OPAMP with PMOS Input Pair
+2-Stage TIA
+3-Stage TIA
+Comparator DC Power (W) Avg. Delay (sec)
+Level Shifter DC Power (W) Ratio Avg. Delay (sec) Delay Balance (sec)
+3.2.2   INSIGHT for Optimization
 The pre-trained INSIGHT is subsequently integrated into a Model-based Batch RL framework for circuit sizing, referred to as INSIGHT-M.
-As shown in the right part of Figure 2, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance target specifications, including previously unseen ones.
-In our experiments, we utilized a Proximal Policy Optimization (PPO)-based RL agent 44, a widely used approach in Reinforcement Learning.
-PPO has been extensively utilized in analog design optimization, as in  45, 24.
-PPO operates by executing multiple trajectory rollouts in parallel, followed by their subsequent aggregation to update the policy under certain constraints.
-Although PPO’s use of multiple trajectories under safe constraints enhances the robustness and safety of the learning process, it tends to compromise sample efficiency.
-The INSIGHT-M framework addresses this drawback by enabling the RL agents to be pre-trained within the close-to-real synthetic neural simulator environment, INSIGHT, until they achieve satisfactory performance.
+As shown in the right part of Figure 2, upon query, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance targets.
+In our experiments, we use PPO-based RL agents  35 with default hyperparameters, a method extensively used in EDA because of its robustness and effectiveness  36, 17.
+PPO enhances learning robustness and safety by executing multiple parallel trajectory rollouts and aggregating them to update policies under safe constraints but at the expense of poor sample efficiency.
+The INSIGHT-M framework addresses this drawback by training robust RL agents within INSIGHT’s synthetic environment, which closely mimics real-world conditions.
 Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy.
@@
 If the agents perform adequately, their deployment persists.
-If not, new real-world data is acquired to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated.
+If not, newly acquired real-world data combined with efficiently sampled previously collected data is utilized to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated.
 This iterative refinement process optimizes the balance between safety, robustness, and efficiency.
-Central to the efficiency of INSIGHT-M is the training of a robust RL agent within a synthetic environment that closely mimics real-world conditions.
 The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency.
-Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, enhances the agent’s adaptability to real-world environments.
-Furthermore, the framework’s capability for iterative refinement effectively balances exploration and exploitation, achieving high sample efficiency.
+Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, further enhances the agent’s adaptability to real-world environments.
+Please note that other RL algorithms can be employed with INSIGHT.
 4. Experiments
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-(a) (b) (c) (d) (e) (f)
-Figure 3.
-Benchmark Circuits: (a) Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, (b) Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, (c) Two-Stage Transimpedance Amplifier, (d) Three-Stage Transimpedance Amplifier, (e) Comparator, (f) Level Shifter.
 Table 4. Performance Comparison Across Different Circuits: INSIGHT vs Neural Simulator in CRONuS
@@
 Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
-Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¿0.99 0.65 ¿0.99 0.62
-Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¿0.99 0.73 ¿0.99 0.77
-Two-Stage Transimpedance Amplifier ¿0.99 0.81 ¿0.99 0.78
-Three-Stage Transimpedance Amplifier ¿0.99 0.91 ¿0.99 0.88
-Comparator 0.70 0.60 0.97 0.61
-Level Shifter ¿0.99 ¡0.00 0.91 0.07
+Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 0.9901 0.6578 0.9960 0.6240
+Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 0.9937 0.7270 0.9973 0.7872
+Two-Stage Transimpedance Amplifier 0.9982 0.8020 0.9933 0.7784
+Three-Stage Transimpedance Amplifier 0.9997 0.9290 0.9996 0.8894
+Comparator 0.7231 0.5537 0.998 0.6535
+Level Shifter 0.9905 -0.1554 0.9265 0.1252
 Circuits (45nm) MSE
@@
 Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
-Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¡0.001 0.01 ¡0.001 0.01
-Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¡0.001 0.05 ¡0.001 0.04
-Two-Stage Transimpedance Amplifier ¡0.001 0.05 ¡0.001 0.05
-Three-Stage Transimpedance Amplifier ¡0.001 0.01 ¡0.001 0.01
-Comparator 0.010 0.01 ¡0.001 0.01
-Level Shifter ¡0.001 0.03 ¡0.001 0.05
+Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 0.00053 0.02054 0.00054 0.01924
+Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 0.00023 0.04267 0.00037 0.04267
+Two-Stage Transimpedance Amplifier 0.00040 0.04335 0.00024 0.04475
+Three-Stage Transimpedance Amplifier 0.00012 0.01688 0.00018 0.01727
+Comparator 0.00666 0.01856 0.00160 0.01856
+Level Shifter 0.00160 0.04436 0.00047 0.05950
 Table 5. INSIGHT Performance Across Different Technologies
 Technology 45nm 130nm 180nm
-Circuit Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
-Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
-3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
-Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
-3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
-Two-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
-3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
-Three-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
-3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
+Circuits Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
+Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 0.9901 0.00053 1500-500 0.9988 0.00027 1500-500 0.9961 0.00046
+3000-1000 0.9960 0.00054 3000-1000 0.9990 0.00082 3000-1000 0.9995 0.00094
+Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1500-500 0.9937 0.00023 1500-500 0.9982 0.00041 1500-500 0.9971 0.00011
+3000-1000 0.9973 0.00037 3000-1000 0.9951 0.00010 3000-1000 0.9961 0.00019
+Two-Stage Transimpedance Amplifier 1500-500 0.9982 0.00040 1500-500 0.9975 0.00065 1500-500 0.9901 0.00050
+3000-1000 0.9933 0.00024 3000-1000 0.9991 0.00023 3000-1000 0.9980 0.00047
+Three-Stage Transimpedance Amplifier 1500-500 0.9997 0.00012 1500-500 0.9956 0.00085 1500-500 0.9969 0.00011
+3000-1000 0.9996 0.00018 3000-1000 0.9977 0.00011 3000-1000 0.9955 0.00019
 Table 6. Performance Comparison of INSIGHT-M (with PPO) vs AutoCkt for 45nm case
-Algorithm INSIGHT-M (with PPO) AutoCkt
-Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1,500 steps 14,100 steps
-Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1,500 steps 9,300 steps
-Two-Stage Transimpedance Amplifier 300 steps 11,100 steps
-Three-Stage Transimpedance Amplifier 600 steps 8,700 steps
-Comparator 1,200 steps 8,400 steps
-4.1. Benchmark Tasks and Algorithms
-In this paper, we utilized a GPU machine with x8 NVIDIA P100 GPUs under Intel(R) Xeon(R) CPU E5-2698 V4 2.220GHz for training and inference.
-We used the following benchmark circuits to evaluate INSIGHT and to validate our approach: Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, Two-Stage Transimpedance Amplifier, Three-Stage Transimpedance Amplifier, Comparator, and Level Shifter.
+Algorithm INSIGHT-M (with PPO) AutoCkt (PPO)
+Circuits Pre-training Data Real-time Simulations Real-time Simulations
+Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1,500 $2\pm 0.18$ steps 14,100 steps
+Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1,500 $2\pm 0.18$ steps 9,300 steps
+Two-Stage Transimpedance Amplifier 300 $15\pm 5$ steps 11,100 steps
+Three-Stage Transimpedance Amplifier 600 $15\pm 5$ steps 8,700 steps
+Comparator 1,200 $15\pm 5$ steps 8,400 steps
+4.1. Benchmark Circuits and Technologies
+We trained and evaluated INSIGHT on a machine with eight NVIDIA P100 GPUs and an Intel Xeon E5-2698 V4 CPU, using the following benchmark circuits to validate our approach: 2-Stage Miller-Compensated OPAMP with NMOS Input Pair, 2-Stage Miller-Compensated OPAMP with PMOS Input Pair, 2-Stage TIA (Transimpedance Amplifier), 3-Stage TIA, Comparator, and Level Shifter.
 Figure 3 illustrates the detailed topology of the benchmark circuits.
-Table 6 presents the performance specifications for each circuit we focussed on for this paper.
-We have validated our concept using BSIM 45nm, Skywater 130nm, BSIM 180nm, and FreePDK 45nm technology nodes.
+Table 3 lists the performance metrics for each circuit we focussed on in this paper.
+We have validated our concept using BSIM 45nm, Skywater 130nm, BSIM 180nm, and FreePDK 45nm technologies.
 4.2. INSIGHT Performance
 4.2.1   Prediction Accuracy
-Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  27, which serves as the baseline for our comparison with INSIGHT.
-Table 3 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on the BSIM 45nm technology node.
-As demonstrated in the table, our analysis for the comparator and level shifter cases involved a combination of DC and simulation-costly transient specifications.
-Specifically for the level shifter, leveraging its autoregressive capabilities, INSIGHT enhances the predictability of simulation-intensive transient specifications such as Avg.
-Delay and Delay Balance by efficiently utilizing economical metric information like DC Power and Ratio.
-The table further details INSIGHT’s performance across various training and testing data sizes, indicating robust performance for dataset sizes close to 1500-1800.
+Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  25, which serves as the baseline for our comparison with INSIGHT.
+We also validate INSIGHT across various types of performance metrics.
+As illustrated in Table 3, different combinations of economic metrics and simulation-intensive transient specifications have been considered for the comparator and level shifter circuits.
+Table 4 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on 45nm technologies.
+The table further details INSIGHT’s performance across various training dataset sizes, indicating robust performance for dataset sizes $\sim$ 1500-1800.
 Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times.
 In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases.
-Table 4 validates INSIGHT’s effectiveness across 45nm, 130nm, and 180nm technology nodes, reinforcing its universal applicability.
-Our experiments showed that simple transfer learning or knowledge-based domain adaptation techniques can reduce initial dataset requirements.
+INSIGHT also demonstrates high prediction accuracies for simulation-intensive transient specifications given economic metric information.
+For example, using just DC Power information significantly improves INSIGHT’s prediction for Avg.
+Delay in comparators.
+Similarly, using just DC Power and Ratio information significantly improves INSIGHT’s prediction accuracy for Avg.
+Delay and Delay Balance.
+Table 5 presents INSIGHT’s prediction accuracy scores across 45nm, 130nm, and 180nm technologies when trained from scratch, using different dataset sizes.
+By employing simple transfer learning, we were able to reduce train dataset requirements successfully.
 4.2.2   Inference Speed
@@
 4.3. INSIGHT-M Performance
-We evaluate the performance of INSIGHT-M against AutoCkt 24, an established online learning-based sizing algorithm based on Proximal Policy Optimization (PPO), using default hyperparameter settings.
-Table 5 highlights the performance comparison.
-Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance.
-For these results, the INSIGHT model used within INSIGHT-M was pre-trained on a minimal initial dataset of 300 samples, showcasing the efficacy of our neural simulator INSIGHT in optimization frameworks.
-Table 6. Performance Specifications for each circuit
-Circuit Performance Specifications Width (Capacitor) Range ($\mu$m ($pF$))
-Two-Stage Miller-Compensated OPAMP with NMOS Input Pair $I_{Q}$ (mA) DC Gain (dB) UGBW (MHz) Phase Margin (degrees) [1.0e-1, 3.0e1]
-Two-Stage Miller-Compensated OPAMP with PMOS Input Pair [1.0e-1, 3.0e1]
-Two-Stage Transimpedance Amplifier [1.0e-1, 1.0e2]
-Three-Stage Transimpedance Amplifier [1.0e-2, 3.0e1]
-Comparator DC Power (W) Avg. Delay (sec) [1.0e-1, 3.0e1]
-Level Shifter DC Power (W) Ratio Avg. Delay (sec) Delay Balance (sec) [4.6e-1, 1.0e2]
+We evaluate the performance of INSIGHT-M against AutoCkt 17, an established online learning-based sizing algorithm based on PPO, using default hyperparameter settings.
+Table 6 presents the performance comparison for three different circuit types, OPAMP, TIA, and Comparator, on 45nm technology for three different test cases, each test case corresponding to a fixed target specification, randomly chosen from a set of designer-specified performance specifications.
+Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance (sample efficiency).
+The table also depicts the initial dataset size used to pre-train INSIGHT for each case.
+Without pre-training, the total number of steps (real-time simulator interactions) required falls within this mentioned pre-training dataset size for each case.
+As the table demonstrates, pre-trained INSIGHT enables INSIGHT-M to achieve convergence at only ¡ 20 real-time simulations across circuits for the given RL algorithm, achieving over 100-1000X improvement in sample efficiency over existing sizing methods.
 5. Conclusion
-This paper proposes INSIGHT, a novel, GPU-powered, technology-independent, effective universal neural simulator in the analog front-end automation loop.
-Leveraging self-attention to understand underlying inherent parameter-performance interdependencies, INSIGHT accurately predicts performance across analog circuits and their corresponding metrics over various technology nodes, achieving inference times within microseconds.
-Notably, its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information.
-Due to its low cost and high fidelity, INSIGHT serves as an effective substitute for standard simulators in analog front-end optimization frameworks, facilitating effective and efficient design space exploration and improved sample efficiency.
-Our Model-Based Batch RL framework, INSIGHT-M, which utilizes INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits.
-We rigorously validated our approach on diverse circuit types across technology nodes.
-To our knowledge, this marks the first application of autoregressive transformers in analog front-end design automation.
-We plan to extend this technology to address more complex analog design challenges, including mismatch analysis and design migration.
-We aim to enhance our adaptation techniques to mitigate synthetic to real-world adaptation errors, thereby strengthening the robustness and reliability of our models under dynamic conditions.
-This innovative application of autoregressive transformers opens new avenues for leveraging advanced NLP techniques, including grammatical models and language-based inversion techniques in circuit modeling and optimization, inspiring further innovation in the field.
+This paper introduces INSIGHT, a novel, GPU-powered, technology-agnostic, effective universal neural simulator as a low-cost and high-fidelity substitute for standard circuit simulators in the analog design automation loop.
+INSIGHT facilitates effective and efficient design space exploration for improved sample efficiency.
+Leveraging self-attention to understand underlying inherent parameter-performance interdependencies within a given topology, INSIGHT accurately predicts performance at microsecond inference times across diverse circuit types and technologies.
+INSIGHT also enables accurate prediction of simulation-costly critical transient specifications using less expensive performance metric information.
+Our Model-Based Batch RL framework, INSIGHT-M, which utilizes INSIGHT for analog sizing, only requires ¡ 20 real-time simulations with 100-1000x lower simulation costs across circuits, showcasing the effectiveness of INSIGHT.
+To the best of our knowledge, this marks the first successful application of a universal neural simulator framework leveraging autoregressive Transformers for analog design automation.
+We also plan to extend this technology to address more complex analog design challenges, including mismatch analysis and technology design migration.

### v3

@@
 Various knowledge-based and optimization-based approaches, including equation-based and simulation-based methods, have been proposed for analog design automation.
-While knowledge-based methods, such as  1, 2, 3, 4, are time-consuming and lack scalability, equation-based methods, like  5, 6, 7, 8, are also highly inaccurate on modern technologies.
+While knowledge-based methods, such as  1, 2, 3, are time-consuming and lack scalability, equation-based methods, like  4, 5, 6, 7, are also highly inaccurate on modern technologies.
 Conversely, simulation-based methods, while effective, depend on costly real-time simulations.
-Among these, evolutionary algorithms like particle swarm optimization  9, advanced differential evolution  10, and genetic algorithms  11, 12 require extensive SPICE simulations (sample inefficient).
-Deep Reinforcement Learning (RL) methods like  13, 14, 15, 16, 17, 18, 19, 20 while robust, also involve frequent simulator interactions.
-Meanwhile, Bayesian Optimization  21, 22 suffers from scalability issues due to computational demands associated with model updates.
-Modern algorithms  23, 24, 25 use black-box optimization and model-based RL to enhance sample efficiency, yet they still require numerous costly simulations.
+Among these, evolutionary algorithms like particle swarm optimization  8, advanced differential evolution  9, and genetic algorithms  10, 11 require extensive SPICE simulations (sample inefficient).
+Deep Reinforcement Learning (RL) methods like  12, 13, 14, 15, 16 while robust, also involve frequent simulator interactions.
+Meanwhile, Bayesian Optimization  17, 18 suffers from scalability issues due to computational demands associated with model updates.
+Modern algorithms  19, 20, 21 use black-box optimization and model-based RL to enhance sample efficiency yet require numerous costly real-time simulations.
 Given the vast and complex design space and intricate trade-offs among performance metrics, repeated compute-intensive simulations within the optimization loops become inherent for effective and efficient design space exploration and exploitation to ensure optimal circuit performance.
 This heavy reliance on expensive real-time simulations motivates us to find an efficient and effective surrogate for the standard SPICE simulator to expedite analog design optimization tasks.
-The complexity of parameter-performance relationships in analog circuits renders simple DNNs inadequate, necessitating bigger, more complex models and extensive, costly EDA data while still resulting in suboptimal accuracies 25.
+The complexity of parameter-performance relationships in analog circuits renders simple DNNs inadequate, necessitating bigger, more complex models and extensive, costly EDA data while still resulting in suboptimal accuracies 21.
 Given the high interdependence of performance metrics inherent to analog circuits, we propose to formulate the analog circuit performance prediction problem as an autoregressive generation task, i.e., predicting each remaining simulation-costly performance specification by leveraging provided design parameters and partial performance information (provided or previously-generated) in a sequential manner.
-The autoregressive problem setup enforces the model to automatically reason and capture the inherent interdependency in the challenging analog performance prediction task, easing the model learning complexity, therefore, providing much better prediction quality than vanilla DNN setup 25.
+The autoregressive problem setup enforces the model to automatically reason and capture the inherent interdependency in the challenging analog performance prediction task, easing the model learning complexity, therefore, providing much better prediction quality than vanilla DNN setup 21.
 In this paper, we introduce INSIGHT, a novel, technology-agnostic, universal neural circuit simulator that accurately predicts performance on previously unseen parameters across various circuits and technologies autoregressively, consistently demonstrating R2 scores as high as 0.99 and minimal test losses at microsecond inference times.
@@
 In analog circuits, performance metrics and parameters are highly interdependent.
-By leveraging these interdependencies, we can build effective prediction models.
+By leveraging these interdependencies, we can build effective and data-efficient prediction models.
 For instance, fundamental metrics such as Quiescent Current ($I_{Q}$) and DC Gain, when combined with parameter information, can significantly enhance predictions for other key specifications, like unity-gain bandwidth (UGBW) and phase margin (PM).
@@
 Thus, leveraging less expensive performance metric information can significantly enhance prediction accuracies for critical, simulation-costly, transient specifications such as slew rate and settling time.
-This insight motivates the adoption of autoregressive models for precise performance prediction.
 Furthermore, given the transferable nature of circuit physics across technologies for a given topology, understanding performance interdependencies within one node provides valuable insights applicable to other nodes.
-These insights, combined with the adaptability and transferability of Transformer architectures, motivate the adoption of autoregressive Transformers to build a universal neural simulator that would work efficiently across technologies.
+These insights, combined with the adaptability and transferability of Transformer architectures, motivate the adoption of autoregressive Transformers to build a universal neural simulator that effectively predicts performance across technologies.
 2.2. Transformer Architectures
@@
 These weights highlight the most relevant token information for predicting subsequent tokens and are derived from the dot product of input vectors.
-Unlike traditional RNNs 26, Transformers process data in parallel, enhancing efficiency and scalability.
-While the original Transformer architecture  27 consists of stacked layers of encoders and decoders, decoder-only models like GPT 28 have recently demonstrated notable success in sequence data generation.
+Unlike traditional RNNs 22, Transformers process data in parallel, enhancing efficiency and scalability.
+While the original Transformer architecture  23 consists of stacked layers of encoders and decoders, decoder-only models like GPT 24 have recently demonstrated notable success in sequence data generation.
 Decoder-only Transformers employ solely the decoder component, composed of multi-head self-attention and position-wise feed-forward networks.
 2.3. Figure of Merit
-Given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  23, 29, 24, to evaluate quality of a design.
+Given an objective performance metric $f_{0}(\mathbf{x})$ and $m$ constraint metrics $\{f_{i}(\mathbf{x})\leq 0\mid i=1,\ldots,m\}$, we adopt the Figure of Merit (FoM) definition given in  19, 25, 20, to evaluate quality of a design.
 This is of the following form:
@@
 (2) $\begin{split}P(y_{1},y_{2},\ldots,y_{M}\mid x_{1},x_{2},\ldots,x_{N})=\\ \prod_{i=1}^{M}P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})\end{split}$
-Each term $P(y_{i}|x_{1},x_{2},\dots,x_{N},)$, which quantifies the likelihood of observing the token $y_{i}$, given the design parameters and the predicted performance specifications.
 This objective is equivalent to minimizing the Negative Log Likelihood (NLL), expressed as:
 (3) $\text{NLL}=-\sum_{i=1}^{M}\log P(y_{i}\mid x_{1},x_{2},\ldots,x_{N},y_{1},y_{2},\ldots,y_{i-1})$
-This formulation connects to the contextual approach in autoregressive sequence generation tasks involving decoder-based Transformers like GPT 28 and motivates us to use a decoder-style Transformer model for INSIGHT.
-Figure 1 presents a conceptual diagram illustrating how INSIGHT implements autoregressive modeling to predict performance metrics.
-We utilize the self-attention mechanism to enable the model to adaptively focus on different parts of the input parameter sequence ($x_{1}$ to $x_{N}$) combined with the previously determined performance outputs ($y_{1}$ to $y_{i-1}$), depending on what is most relevant for predicting each specific output metric $y_{i}$.
-Positional encoding is also incorporated to retain the sequence order, crucial for maintaining context and coherence.
-The autoregressive feature ensures the model learns to utilize dependencies of one performance metric on another, along with the parameter information, to predict the performance metrics accurately.
+Each term, $P(y_{i}|x_{1},x_{2},\dots,x_{N},y_{1},y_{2},\ldots,y_{i-1})$, quantifies the likelihood of observing the token $y_{i}$, given the design parameters and the predicted performance specifications.
+This formulation connects to the contextual approach in autoregressive sequence generation tasks involving decoder-based Transformers like GPT 24 and motivates us to use a decoder-style Transformer model for INSIGHT.
+Figure 1 presents a conceptual diagram illustrating INSIGHT’s autoregressive sequential inferencing of performance metrics.
+In autoregressive sequence generation, accurate prediction of one metric significantly enhances the prediction accuracy of the subsequent, emphasizing the importance of metric sequence order for effective and data-efficient predictions.
+We employ a greedy strategy, ordering metrics by increasing complexity and dataset needs.
+Specifically, we start with simpler, cost-effective metrics such as DC and AC and progress to more complex ones like transients.
+Also, if a specification is known to be directly derived from others, it is sequenced after the latter to optimize data need and prediction accuracy.
+By avoiding the early use of complex metrics, which demand more data, this approach leverages simpler specifications to enhance the prediction accuracy of more complex ones, ensuring optimal, effective, and data-efficient sequencing.
+Given a performance metric sequence order, INSIGHT utilizes self-attention to adaptively focus on the most relevant segments of the parameter ($x_{1}$ to $x_{N}$) and generated performance output ($y_{1}$ to $y_{i-1}$) sequence, to accurately predict each output metric $y_{i}$.
+Positional encoding is used to preserve sequence order and maintain context and coherence.
 Table 1 presents INSIGHT’s architecture details.
@@
   # of Decoder Layers 3
-  Activation GeLU 30
+  Activation GeLU 26
   Learning rate 0.001
-with cosine annealing scheduler  31
+with cosine annealing scheduler  27
 Table 2. Neural Simulator Architecture in CRONuS
@@
 # of Ensemble Networks 7
-Activation ReLU 32
+Activation ReLU 28
   Learning rate 0.001
-Although our analog performance prediction task is analogous to the sequence generation tasks, we would like to emphasize that we need to predict real values instead of generating a sequence of tokens, such as a word or a position in an image, from a set of finite elements, or generate scalar real values from sequential contexts, as shown in previous Transformer works.
-Generating a sequence of real values based on Transformer architecture is unclear  33, and thus, suitable foundation models for predicting circuit parameters do not exist.
-Therefore, it is essential to evaluate the feasibility of Transformer architecture in generating a sequence of circuit outputs to confirm the potential for constructing circuit foundation models.
-Accordingly, building on the existing decoder-Transformer architecture 28, 34, we show that our proposed architecture accelerates inference and training to demonstrate the feasibility of a unified framework like a foundation model.
-This streamlined approach enhances processing efficiency, reduces data requirements, and makes the model more practical for various applications.
-By focusing on simplicity, we improve performance and usability while maintaining robust predictions.
+Note: While the relationship between performance metrics, relative to design parameters and among themselves, is not inherently time-series, this formulation enables self-attention and autoregressive features to leverage their interdependencies for effective and efficient predictions.
+Although our task is analogous to sequence generation tasks, here we need to predict real values instead of generating a sequence of tokens, such as words or positions in images, from a set of finite elements or generating scalar real values from sequential contexts, as shown in previous Transformer works.
+Generating a sequence of real values based on Transformer architecture is unclear  29, and thus, suitable foundation models for predicting circuit performance metrics do not exist.
+Therefore, evaluating the feasibility of the Transformer architecture in generating a sequence of circuit outputs is crucial to validate its potential for building domain-specific performance prediction models within a universal framework.
+Building on the existing decoder-Transformer architecture 24, 30, our proposed architecture for INSIGHT, combined with our strategy for optimizing performance sequence order, offers a data-efficient and effective universal framework that seamlessly accommodates various circuit types and topologies, is technology-agnostic and automatically adapts to diverse parameter and performance metric lengths.
+By focusing on simplicity, we improve usability and scalability while maintaining robust predictions.
 3.2. INSIGHT for Analog Sizing
@@
 As shown in the right part of Figure 2, upon query, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance targets.
-In our experiments, we use PPO-based RL agents  35 with default hyperparameters, a method extensively used in EDA because of its robustness and effectiveness  36, 17.
+In our experiments, we use PPO-based RL agents  31 with default hyperparameters, a method extensively used in EDA because of its robustness and effectiveness  32, 16.
 PPO enhances learning robustness and safety by executing multiple parallel trajectory rollouts and aggregating them to update policies under safe constraints but at the expense of poor sample efficiency.
@@
 Circuits (45nm) R2 Scores
-Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
+Trainset:Testset Size Range 300:100 - 1,500:500 3,$0$ 9,000:3,000
 Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
@@
 Circuits (45nm) MSE
-Trainset-Testset 300-100$\sim$1,500-500 3,000-1,000$\sim$9,000-3,000
+Trainset:Testset Size Range 300:100 - 1,500:500 3,$0$ 9,000:3,000
 Neural Simulator INSIGHT FC-Ensemble in CRONuS INSIGHT FC-Ensemble in CRONuS
@@
 Level Shifter 0.00160 0.04436 0.00047 0.05950
-Table 5. INSIGHT Performance Across Different Technologies
+Table 5. INSIGHT Performance Across Different Technologies (Training from Scratch)
 Technology 45nm 130nm 180nm
-Circuits Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
+Circuits Train-Test Size R2 MSE Train-Test Size R2 MSE Train-Test Size R2 MSE
 Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 0.9901 0.00053 1500-500 0.9988 0.00027 1500-500 0.9961 0.00046
@@
 4.2.1   Prediction Accuracy
-Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  25, which serves as the baseline for our comparison with INSIGHT.
+Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  21, which serves as the baseline for our comparison with INSIGHT.
 We also validate INSIGHT across various types of performance metrics.
@@
 Table 5 presents INSIGHT’s prediction accuracy scores across 45nm, 130nm, and 180nm technologies when trained from scratch, using different dataset sizes.
-By employing simple transfer learning, we were able to reduce train dataset requirements successfully.
+By employing simple transfer learning, we could further reduce train dataset requirements.
 4.2.2   Inference Speed
@@
 4.3. INSIGHT-M Performance
-We evaluate the performance of INSIGHT-M against AutoCkt 17, an established online learning-based sizing algorithm based on PPO, using default hyperparameter settings.
-Table 6 presents the performance comparison for three different circuit types, OPAMP, TIA, and Comparator, on 45nm technology for three different test cases, each test case corresponding to a fixed target specification, randomly chosen from a set of designer-specified performance specifications.
-Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance (sample efficiency).
-The table also depicts the initial dataset size used to pre-train INSIGHT for each case.
-Without pre-training, the total number of steps (real-time simulator interactions) required falls within this mentioned pre-training dataset size for each case.
+We evaluate the performance of INSIGHT-M against AutoCkt 16, an established online learning-based sizing algorithm based on PPO, using default hyperparameter settings.
+Table 6 presents the performance comparison for different circuits on 45nm technology, each row representing a test case corresponding to a fixed target specification randomly selected from a set of designer-specified performance specifications similar to those used in CRONuS 21 and AutoCkt.
+Performance is measured based on the number of real-time simulations (steps) required, with fewer steps indicating better performance (sample efficiency).
+The performance is evaluated over 30 trials to ensure robust statistical analysis, with INSIGHT-M’s reported as mean $\pm$ standard deviation.The table also depicts the initial dataset size used to pre-train INSIGHT for each case.
+Without pre-training, the framework behaves like a pure Model-based RL setup, with the total number of steps required falling within this mentioned pre-training dataset size for each case.
 As the table demonstrates, pre-trained INSIGHT enables INSIGHT-M to achieve convergence at only ¡ 20 real-time simulations across circuits for the given RL algorithm, achieving over 100-1000X improvement in sample efficiency over existing sizing methods.